OpenStudy (anonymous):

Judy works at a local bowling alley. She currently makes $300 per week after taxes plus $60 extra cash every other week for cleaning the lanes. Judy is trying to save up for a car. She really wants a Bug. She knows that she can buy a used Bug in good condition for between $6,600 and $9,900. The equation below represents this situation where x is the number of weeks worked and P is the price of the car: 1/2(60x) + 300x = P What is the minimum number of weeks that Judy needs to work to buy her Bug? A.)30 weeks B.)20 weeks C.)25 weeks D.)10 weeks

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }(60x)+300x= \frac{ 60x }{ 2 }+300x\] \[\frac{ 60x }{ 2 }=30x\] then you get:\[30x+300x=P\] you can combine the terms and get \[330x\] since it says minimum, you take $6600 as your minimum goal thus you do \[330x=6600\rightarrow \frac{ 330x }{ 330 }=\frac{ 6600 }{ 330 }\] \[x=20\]